OpenStudy (anonymous):

solve: 3/x-2+8/x+3=2 show working please :)

OpenStudy (anonymous):

Do you mean x under denominator like this? 3/(x-2) + 8/(x+3) = 2

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Common denominator is (x +3) (x-2) 3 ( x +3) + 8 (x -2) = 2 (x +3) (x-2) 3x +9 + 8x - 16 = 2 ( x^2 +x -6) 2x^2 + 2x - 12 = 3x +9 + 8x - 16 2x^2 -9x -5 = 0 x = (+9 +/- 11 )/4 = x = 5, x = -1/2

OpenStudy (anonymous):

yayyyy thank you! :D